How would you solve this absolute value equation (the little lines) w+1 (the little lines) equals five
Answer by Alan3354(69443)

|w+1| = 5
---------
Case 1:
w+1 = 5
W = 4
---------
Case 2:
-(w+1) = 5
w+1 = -5
w = -6
--------
That's how I would do it.